you have a chemical reaction between 700 grams of compound A & 300 grams of compound B which produces compound C. You measure the resultant amount of C in grams over 8 trials which are listed below. What range do you estimate for the amount of compound C that will be produced by any single performance of this reaction? You should give a range that you are 95% confident in.

676.3
546.6
456.6
323.9
665.4
546.8
675.3
434.5

amount of c produced = [ ] ± [ ]

please show all steps. thank in advance!
Answer #1

sample std dev ,    s = √(Σ(X- x̅ )²/(n-1) ) =   129.5487
Sample Size ,   n =    8
Sample Mean,    x̅ = ΣX/n =    540.6750

Level of Significance ,    α =    0.05          
degree of freedom=   DF=n-1=   7          
't value='   tα/2=   2.3646   [Excel formula =t.inv(α/2,df) ]      
                  
Standard Error , SE = s/√n =   129.5487   / √   8   =   45.802393
margin of error , E=t*SE =   2.3646   *   45.80239   =   108.305449
                  
confidence interval is                   
Interval Lower Limit = x̅ - E =    540.68   -   108.305449   =   432.3696
Interval Upper Limit = x̅ + E =    540.68   -   108.305449   =   648.9804
95%   confidence interval is (   432.37   < µ <   648.98   )